OpenStudy (anonymous):

Hi! So I'm asked to simplify logs: log(base 3)X+5log(base1/3)X I have no idea how to simplify with different bases. Is it even possible?

OpenStudy (kc_kennylau):

Yes it is possible

OpenStudy (kc_kennylau):

\[\log_ab=\frac{\log b}{\log a}\]

OpenStudy (kc_kennylau):

log(base 3)X+5log(base1/3)X = log(base 3)X+5log(base3)X/log(base3)(1/3)

OpenStudy (kc_kennylau):

You can do it now?

OpenStudy (anonymous):

I thought it would be log X/ log 1/3 if I used that property? Since A=1/3 and B=X

OpenStudy (kc_kennylau):

Well base 10 is just the same as any base, and sorry I should modify my formula:

OpenStudy (kc_kennylau):

\[\log_ab=\frac{\log_nb}{\log_na}\]

OpenStudy (anonymous):

I'm confused as to why it would be 5log(BASE 3)X/log(BASE 3)X like how is it that instead of 10 it would be 3 for the base

OpenStudy (kc_kennylau):

Well I put n=3